OpenStudy (anonymous):

How do I solve the problem a-59/18=4/3(a+1/3)-5/3?

OpenStudy (anonymous):

\[a-\frac{59}{18}=\frac{4}{3}*(a+\frac{1}{3})-\frac{5}{3}\]

hero (hero):

Which is it?

OpenStudy (anonymous):

The first response, a is not above the fraction line

hero (hero):

\(\large a - \frac{59}{18} = \frac{4(a + \frac{1}{3})}{3} - \frac{5}{3}\)

OpenStudy (anonymous):

The answer that is given on the paper states it's 1/2, but I don't know how to come to that conclusion

OpenStudy (anonymous):

Okay, I see what you did with 4/3

hero (hero):

\(\large a - \frac{4(a+\frac{1}{3})}{3} = \frac{59}{18} - \frac{5}{3}\)

hero (hero):

Multiply a by 3/3 and 5/3 by 6/6 to get: \(\large \frac{3a}{3} - \frac{4(a + \frac{1}{3})}{3} = \frac{59}{18} - \frac{30}{18}\) Are you still with me?

OpenStudy (anonymous):

Yes

hero (hero):

Combine fractions to get \(\large\frac{3a - 4(a + \frac{1}{3})}{3} = \frac{29}{18}\)

hero (hero):

Simplify the left numerator to get: \(\large\frac{(3a - 4a - \frac{1}{3})}{3} = \frac{29}{18}\)

OpenStudy (anonymous):

Okay

hero (hero):

Simplify further to get: \(\large\frac{-a - \frac{1}{3}}{3} = \frac{29}{18}\)

hero (hero):

Extract the negative sign to get: \(\large -\frac{a + \frac{1}{3}}{3} = \frac{29}{18}\)

hero (hero):

Cross multiply to get -18(a + 1/3) = 3(29)

hero (hero):

Simplify again to get: -18a - 6 = 87 Add 6 to both sides to get: -18a = 93 Add 18a to both sides; subtract 93 from both sides to get -93 = 18a Divide both sides by 18:

OpenStudy (anonymous):

Okay

hero (hero):

\(a = -\frac{93}{18}\) but that reduces to \(a = -\frac{37}{6} \dot\ \frac{6}{6}\) \(a = -\frac{37}{6} \dot\ 1\)

hero (hero):

Wow your teacher wanted to torture you

hero (hero):

That's the correct answer because I verified it with my calculator. Maybe you posted the problem wrong?

OpenStudy (anonymous):

I know T_T 3 days into freshman year and I get asked to do this xd

hero (hero):

We didn't get 1/2

OpenStudy (anonymous):

Nope, well actually I've gotte nthat same answer too, it could be a fault on the question

hero (hero):

Is there a way you can upload the question from the book?

hero (hero):

Scan it

OpenStudy (anonymous):

It's on a paper from school

hero (hero):

Exactly, so scan it or use the drawing button below.

hero (hero):

Use the drawing button below to write it out exactly how you see it.

OpenStudy (anonymous):

Alright I'll draw it because it takes a while to scan

OpenStudy (anonymous):